a car travels equal distances with successive speeds of 10, 30, 60 and 20 kilometre per hour to go from one point to another point along a straight line its average speed is

Grade:12th pass

1 Answers

Vikas TU
7 years ago
Let x be the distance in each case.
x = 10*t1
and
x = 30*t2
and
x = 60*t3
and
x = 20*t4
and
t = t1 + t2 + t3 + t4
avg. speed = total distance / total time
                = 4x/t => 4x/(x/10 + x/30 + x/60 + x/20) 
                        => 4/((6+2+1+3)/60))
                        => 4*60/12
                        => 20 Km/hr.